OpenStudy (anonymous):

A 24-foot pole is leaning against a wall. It touches the wall 12 feet above the ground. Find the measure of the angle that the pole makes with the ground. Round your answer to the nearest degree.

OpenStudy (anonymous):

sin(θ) = 12/24 = 1/2 θ = arcsin(1/2) θ = 30°